OpenStudy (anonymous):

Solve (2x+5)(4x−3)=0 A. x = 5/2 or x = -3/4 B. x = -5/2 or x = 3/4 C. x = 2/5 or x = 4/3 D. x = -2/5 or x = -4/3

OpenStudy (anonymous):

solve each term for 0 2x+5=0 4x−3=0 Then you will have your two answers.

OpenStudy (anonymous):

Example with different values: 9x+3= 0 subtract 3 from each side to start isolating x = 9x = -3 Then to isolate x, divide both sides by 9 = x = -3/9 simplify/reduce x=-1/3